OpenStudy (anonymous):

what are two unit vectors that are orthogonal to both <1,-1,1> and <0,4,4>

OpenStudy (anonymous):

Cross product..

OpenStudy (anonymous):

Im not getting the right answer for some reason

OpenStudy (anonymous):

Normalize!

OpenStudy (anonymous):

can you show me the steps?

OpenStudy (anonymous):

\[ \frac{v\times w}{||v\times w||} \]

OpenStudy (anonymous):

I don't understand the denominator

OpenStudy (anonymous):

I guess my question would be how do you normalize

OpenStudy (anonymous):

Well if you have a vector v=(x,y,z) then \[ ||v|| := \sqrt{x^2+y^2+z^2} \]

OpenStudy (zarkon):

\[\|<a,b,c>\|=\sqrt{a^2+b^2+c^2}\]

OpenStudy (anonymous):

ok, I got one of them. Would the other orthogonal vector just be B X A ?

OpenStudy (anonymous):

Just take the negative!

OpenStudy (anonymous):

Multiply by -1.

OpenStudy (anonymous):

oh, you're right!

OpenStudy (anonymous):

haha thanks.